# Towards Goal-Oriented Semantic Communications: New Metrics, Framework, and Open Challenges

### v2

@@ Abstract
Since Shannon’s pioneering masterpiece which established the foundation of modern information theory, the design target of communications has long been promising bit-level message reconstruction and achieving shannon capacity. However, this focus neglects the semantics and effectiveness aspects of information. Recently, the development of wireless technologies and the spurt of deep learning (DL) techniques allow us to reclaim the meaning/usefulness aspect in the design of 6G communication paradigms, where goal-oriented communication is becoming a trend. Age of Information (AoI), a well-known metric that captures the significance of information by recording the time elapsed from the generation time slot, has been extended to various variants, such as Value of Information (VoI), Urgency of Information (UoI), Age of Incorrect Information (AoII), and etc. While each of them proposes novel ways to measure the semantics/effectiveness aspect of information, there is not yet a unified framework encompassing all of them. To this end, we propose a novel tensor-based approach, the Goal-oriented Tensor (GoT), to unify them, which also allows more flexible and fine-grained goal characterizations. Following the proposed GoT, we architect a holistic goal-oriented semantic communications framework, in which semantics perception, dissemination, and control-plane decision making collaboratively work towards realizing specific goals. Finally, we outline several open challenges to fulfill the vision of the GoT framework.
@@ I-A Beyond Bit-Oriented Transmission
The modern prosperity of communication technologies is primarily standing on the cornerstone built by Shannon’s Information Theory in 1948 [1]. Since its inception, a primary goal in the field of communication has been to approach Shannon’s theoretical capacity for transmission. Today, the advanced 5G technology has nearly achieved this goal. Technically, advances like massive multiple-input multiple-output (MIMO), and millimeter wave communications have enabled a plethora of services with ultra reliability, low latency, and high data rate. At the physical layer, the combination of Low-Density Parity-Check (LDPC) codes and Polar codes also represents a breakthrough in realizing low-complexity near-Shannon-limit coding-decoding pairs. Such technologies aim at building perfect bit-by-bit information pipelines between transceivers, which follow Shannon’s legacy precisely.
@@ I-B Overview of Semantic Communications
There are two primary avenues that lead the development of semantic communications. The first avenue applies DL-learning techniques to extract semantic features and filter semantics-irrelevant redundancy within a specific packet, thus enhancing communication efficiency. The second avenue interprets semantics as its etymological meaning, the significance of information (also known as priority of information or significance of information).
I-B1 Avenue 1: Semantics as Meaning
@@
I-B2 Avenue 2: Semantics as Significance
The second avenue interprets semantics as “significance”, “priority”, or “significance[7], which is also our focus in this article. From this perspective, intelligent networked systems can identify high-priority messages and adaptively allocate more resources. Notably, the determination of priority deviates from traditional, probability-centric entropy measures. Instead, it leans on the “significance” attributed to the goal-oriented usefulness of message. As noted in [8], “Imagine two equally rare events, occurring with very low probability, one of which carries a major safety risk while the other is just a peculiarity. Although they provide the same high amount of information, the information conveyed by the first event is evidently of higher significance.”, it holds great potential to design a semantics/goal-aware relevant filter to slim down the information pipeline, where only packets that carry relevant meaning for accomplishing the goal are prioritized for transmission. Since the significance relates to the goal, the significance interpretation also bridges this avenue to the effectiveness level. In the second avenue, the core challenge is to design metrics that could characterize the significance property. Age of Information (AoI), defined as the time elapsed since the generation of the most recent information update at the destination, is one typical metric that captures the significance property. In the AoI-oriented system, fresher messages are deemed more relevant to the end-user, and thus prioritized for transmission. However, AoI does not perceive the content of the packet, nor the task at the receiver. To address this issue, improved non-linear variants of AoI, such as VoI [9], AoII [10], UoI [11], Cost of Actuation Error [12], and etc., have been developed. However, there is not yet a unified framework that encompass all these metrics. Moreover, the existing metrics do not directly characterize the desired goal. These issues motivate us to propose a unified, flexible, and directly goal-oriented metric that could assist in accurately capturing the underlying goal, facilitating decision-making, and ultimately improving effectiveness.
@@ I-C Towards Goal-Oriented Semantic Communications
This article addresses the limitations in the second avenue of semantic communication, focusing on the challenge of designing metrics for directly measuring information significance at the effectiveness/goal-oriented level. First, this article comprehensively reviews the existing metrics that capture the significance property in Section II. In this way, we reveal how these metrics manifest the effectiveness level. Then, upon examining the inherent relationships among these metrics, we are inspired to unify them in a more cohesive framework. This leads to the proposed Goal-oriented Tensor (GoT). We visualize examples to demonstrate how the GoT metric reduces to existing metrics that captures significance of information and how it enables more fine-grained and flexible goal characterizations. Consequently, abstract goals could be defined and quantified via GoT.
Furthermore, this article illustrates that information traverses a perception-actuation life cycle to achieve a certain goal. Within this life cycle, information initiates with its generation, goes through semantics extraction (the meaning-aware approach will enhance the effectiveness), coding & modulation, data dissemination, and ends with its transformations into effective decisions (or control) towards goals. In this way, the significance of information is not solely tied to the packet’s inherent meaning, but the ultimate usage of the meaning for achieving particular goals. Incorporating the life cycle, this article then envisions a holistic framework of goal-oriented semantic communications, whereby semantics perception, dissemination, and control-plane decisions are orchestrated in harmony with a shared goal characterized by GoT. We consider an easy-to-follow case study, a real-time wireless fire monitor and rescue system, to illustrate the universality of our proposed framework. We illustrate how the GoT describes the goal in real scenarios. The idea of goal and semantics-aware filter (an open challenge proposed in [7] and [8]) is also explored and verified in this article.
II Capturing the significance of Information: Critical Metrics
Recent studies have explored various metrics to assess the significance of information. This section reviews these metrics, with an emphasis on how these metrics interconnect with the effectiveness of communications.
@@ II-A Age of Information (AoI)
Age of Information (AoI), proposed in 2012 [13], is the first concrete and quantitative metric to characterize the significance of information. Intuitively known as freshness of information, AoI was designed to fulfill the freshness requirements for the proliferated machine-type communications, such as intelligent vehicle networks, industrial internet of things, and many other ultra reliable -supported applications. Central to optimizations on AoI is a subtle yet vital consensus: fresher messages contain more valuable information. This consensus is rooted in the understanding that fresh information ensures more informed and goal-oriented decision-making.
@@ II-C Mean Square Error (MSE)
MSE is a well-known metric that characterizes the accuracy aspect of information. MSE is defined as the long term average squared error (SE) between transceivers’ statuses, which is used for timely reconstruction-oriented communications. In the MSE-oriented design, a status at the source is deemed significant if it considerably deviates from the estimated status at the monitor. When the sample policy is content-agnostic, the MSE can be expressed as a nonlinear function of AoI [14], making it a specialized form of VoI. MSE emphasizes the impact of the severity of E2E status mismatch on the effectiveness.
@@ II-E Age of Incorrect Information (AoII)
AoII is a new metric that orchestrates the strengths of MSE and AoS to measure significance property [10]. As shown in Table I, AoII introduces three key innovations: First, AoII integrates AoS with MSE by multiplying the variants of them together, which incorporates both the duration and severity of E2E mismatch. Second, AoII incorporates a nonlinear penalty function predicated on AoS, drawing parallels to the penalty function in the Value of Information (VoI) relative to AoI. Third, AoII introduces a generalized error gap function, broadening the traditional Euclidean-distance-based MSE depictions. In this regard, AoII addresses that both severity and duration of E2E statues mismatch affect the effectiveness.
@@ II-G Urgency of Information (UoI)
While previous metrics focus solely on the significance of the source, practical scenarios demonstrate that the surrounding environment also influences information significance. For example, a self-driving car in complex situations like traffic jams would require more frequent status updates for safety, unlike when it’s on a clear highway. Addressing this, UoI is the first metric to link environment factors with information significance [11]. Particularly, UoI introduces an environment-aware weighted coefficient based on the environment urgency level. The final UoI formula is a product of this coefficient and the error gap function, as detailed in Table I.
@@ III One More Step Forward: the Goal-oriented Tensor
The overview and discussions in Secion II leads to two key insights: the existing metrics’ assessment of effectiveness is indirect and they lack a unified approach. In this section, we introduce a unified metric to capture the significance of information in a goal-oriented manner and seek to bring coherence to these various metrics.
@@ III-A Existing Metrics: Inherent Connection
From Section II and Table I, the existing metrics share two fundamental components: $i$) a content-aware error cost function $g(X(t),\hat{X}\left(t\right))$ and $ii$) a content-independent weighted coefficient $\Phi(t)$. The former one describes the real-time E2E distance (like MSE or other error gap function) or mismatch cost (like Cost of Actuation Error). The latter one exerts a multiplicative effect on the error cost function. Thus, the inherent connection among the existing metrics is that they are all involve a triple-tuple consisting of $X\left(t\right)$, $\hat{X}\left(t\right)$, and $\Phi\left(t\right)$. This connection motivates us to propose a tensor-based approach to unify the existing metrics.
@@ III-C Constructing Generalized Goal-Oriented Tensor
The values in the generalized GoT manifest the cost of E2E statuses mismatch under a specific context status. Here a detailed method to tailor a GoT based on the goal is given.
@@
[…]
$\bullet$ Step 4: Evaluating the specific cost. There are three types of costs: $i$) the status inherent cost $\mathcal{C}_{1}(X(t),\Phi(t))$, which represents the cost per time slot under $X\left(t\right)$ and $\Phi\left(t\right)$, for instance, $\mathcal{C}_{1}(\text{No Fire},\text{Sunny})=0$ represents zero cost per time slot when there is no fire in a sunny day; $ii$) the decision gain $\mathcal{C}_{2}(\pi(\hat{X}(t)))$ is the cost reduction from an effective action $\pi(\hat{X}\left(t\right))$, for instance, $\mathcal{C}_{2}(\text{Rescue})=5$ represents that the rescue action can reduces 5 unit cost per time slot; $iii$) the action cost $\mathcal{C}_{3}(\pi(\hat{X}(t)))$ corresponds to the resource overhead due to the action $\pi(\hat{X}(t))$, for instance, $\mathcal{C}_{3}(\text{Rescue})=3$ represents that the rescue action incur 3 unit cost per time slot.
$\bullet$ Step 5: Calculating the GoT. The tensor value, given a specific triple-tuple $\langle X(t),\hat{X}(t),\Phi(t)\rangle$ and a determined decision mapping approach $\pi(\cdot)$, is calculated by
$[\mathcal{C}_{1}(X(t),\Phi(t))-\mathcal{C}_{2}(\pi(\hat{X}(t)))]^{+}+\mathcal{C}_{3}(\pi(\hat{X}(t))),$ (1)
[…]
To conclude, the tensor value within GoT is determined by both the E2E semantic mismatch and the decision strategy [^1] Decision strategy can serve as an extended dimension of the GoT. It will change the value of the 3-dimensional GoT., representing the real-world cost per time slot due to these mismatches. Through GoT optimization, we enhance decision-making and mitigate significant E2E semantic mismatch impact. Therefore, the long-term real-world cost is effectively reduced, highlighting GoT’s direct focus on effectiveness.
@@ IV A Holistic Framework for Goal-Oriented Semantic Communication
This section introduces a comprehensive goal-oriented semantic framework, depicted in Fig. 2, with the GoT acting as the central metric. The modules are elaborated as follows:
Semantic Quantization. Status update packets are generated when the sensors collect continuous raw data. The types of metadata are various, e.g., a piece of video/speech, a process of temperature variations, a fragment of moving track, or even their multi-modal combinations. Raw data are fed into a semantic quantizer for discrete semantic segments, producing semantic representation $X(t)$ and context representation $\Phi(t)$. The focus of the quantizer is to convert continuous, non-meaningful metadata into discrete semantic pieces, and thus we call this process as semantic quantization. This method of discretization offers dual benefits: it reduces the data volume per packet, thereby speeding up further processing and saving communication resources; and it enables the identification of semantics priorities in terms of goals, facilitating the sparse semantics-aware sampling design.
@@
Goal-oriented Channel Coding. Traditionally, channel coding strategies aim for perfect reconstructions and lower average decoding error rate. However, this conventional approach treats all end-to-end (E2E) errors equivalently, disregarding the varying consequences associated with different types of errors. To address this issue, we can introduce goal-oriented channel coding to minimize the average cost associated with errors. In this context, E2E errors with potentially severe impacts are allocated larger code distances, thereby diminishing the likelihood of such errors, vice versa. This method ensures that more critical errors are less probable, thus enhancing system goal-achieving utility.

### v3

@@ Abstract
Since Shannon’s pioneering masterpiece which established the foundation of modern information theory, the design target of communications has long been promising bit-level message reconstruction and achieving shannon capacity, which neglects the semantics and effectiveness aspects of information. Nevertheless, the recent development of wireless technologies and the spurt of deep learning (DL) techniques allow us to reclaim the meaning/usefulness aspect in the design of future 6G communication paradigms, where goal-oriented communication is becoming a trend. Age of Information (AoI), a well-known metric that captures the significance of information by recording the time elapsed from the generation time slot, has been extended to various variants, such as Value of Information (VoI), Urgency of Information (UoI), Age of Incorrect Information (AoII), and etc. While each of them proposes novel ways to measure the semantics/effectiveness aspect of information, there is not yet a unified framework encompassing all of them. To this end, we propose a novel tensor-based approach, the Goal-oriented Tensor (GoT), to unify them, which also allows more flexible and fine-grained goal characterizations. Following the proposed GoT, we architect a holistic goal-oriented semantic communications framework, in which semantics perception, dissemination, and control-plane decision making collaboratively work towards realizing specific goals. Finally, we outline several open challenges to fulfill the vision of the GoT framework.
@@ I Introduction
The modern prosperity of communication technologies is primarily standing on the cornerstone built by Shannon’s Classical Information Theory in 1948 [1]. Since its inception, a primary goal in the field of communication has been to approach Shannon’s theoretical capacity for transmission. Today, the advanced 5G technology has nearly achieved this goal. Technically, advances like massive multiple-input multiple-output (MIMO), and millimeter wave communications have enabled a plethora of services with ultra reliability, low latency, and high data rate. At the physical layer, the combination of Low-Density Parity-Check (LDPC) codes and Polar codes also represents a breakthrough in realizing low-complexity near-Shannon-limit coding-decoding pairs. Such technologies aim at building perfect bit-by-bit information pipelines between transceivers, which follow Shannon’s legacy precisely.
@@ I-B Overview of Semantic Communications
There are two primary avenues that lead the development of semantic communications. The first avenue applies DL-learning techniques to extract semantic features and filter semantics-irrelevant redundancy within a specific packet, thus enhancing communication efficiency. The second avenue interprets semantics as its etymological meaning, the significance of information (also known as priority of information or relevance of information).
@@
The second avenue interprets semantics as “significance”, “priority”, or “relevance[7], which is also our focus in this article. From this perspective, intelligent networked systems can identify high-priority messages and adaptively allocate more resources. Notably, the determination of priority deviates from traditional, probability-centric entropy measures. Instead, it leans on the “relevance” attributed to the goal-oriented usefulness of message. As noted in [8], “Imagine two equally rare events, occurring with very low probability, one of which carries a major safety risk while the other is just a peculiarity. Although they provide the same high amount of information, the information conveyed by the first event is evidently of higher significance.”, it holds great potential to design a semantics/goal-aware relevant filter to slim down the information pipeline, where only packets that carry relevant meaning for accomplishing the goal are prioritized for transmission. Since the relevance relates to the goal, the relevance interpretation also bridges this avenue to the effectiveness level, motivating us to expand upon goal-oriented semantic communications from this viewpoint.
In the second avenue, the core challenge is to design metrics that could characterize the relevance property. Age of Information (AoI), defined as the time elapsed since the generation of the most recent information update at the destination, is one typical metric that captures the relevance property. In the AoI-oriented system, fresher messages are deemed more relevant to the end-user, and thus prioritized for transmission. However, AoI does not perceive the content of the packet, nor the task at the receiver. To address this issue, improved non-linear variants of AoI, such as VoI [9], AoII [10], UoI [11], Cost of Actuation Error [12], and etc., have been developed. However, there is not yet a unified framework that encompass all these metrics. Moreover, the existing metrics do not directly characterize the desired goal. These issues motivate us to propose a unified, flexible, and directly goal-oriented metric that could assist in accurately capturing the underlying goal, facilitating decision-making, and ultimately improving effectiveness.
@@ I-C Towards Goal-Oriented Semantic Communications
Following the second avenue, this article addresses the challenge of metrics design. First, this article comprehensively reviews the existing metrics that capture the relevance property in Section II. In this way, we reveal how these metrics manifest the effectiveness level. Then, upon examining the inherent relationships among these metrics, we are inspired to unify them in a more cohesive framework. This leads to the proposed Goal-oriented Tensor (GoT). We visualize examples to demonstrate how the GoT metric reduces to existing metrics that captures relevance and how it enables more fine-grained and flexible goal characterizations. Consequently, abstract goals could be defined and quantified via GoT.
Furthermore, this article illustrates that information traverses a perception-actuation life cycle to achieve a certain goal. Within this life cycle, information initiates with its generation, goes through semantics extraction (the meaning-aware approach will enhance the effectiveness), coding & modulation, data dissemination, and ends with its transformations into effective decisions (or control) towards goals. In this way, the relevance of information is not solely tied to the packet’s inherent meaning, but the ultimate usage of the meaning for achieving particular goals.
Incorporating the life cycle, this article then envisions a holistic framework of goal-oriented semantic communications, whereby semantics perception, dissemination, and control-plane decisions are orchestrated in harmony with a shared goal characterized by GoT. We consider an easy-to-follow case study, a real-time wireless fire monitor and rescue system, to illustrate the universality of our proposed framework. Through this example, we illustrate how the GoT describes the goal in real scenarios. The idea of goal and semantics-aware filter (an open challenge proposed in [7] and [8]) is also initially achieved in this article.
II Capturing the Relevance of Information: Critical Metrics
Recently, many metrics have been investigated to characterize the relevance property of information. This section provides an overview of these metrics and discusses their influence on effectiveness.
@@ II-A Age of Information (AoI)
Age of Information (AoI), proposed in 2012 [13], is the first concrete and quantitative metric to characterize the relevance of information. Intuitively known as freshness of information, AoI was designed to fulfill the freshness requirements for the proliferated machine-type communications, such as intelligent vehicle networks, industrial internet of things, and etc. Central to optimizations on AoI is a subtle yet vital consensus: fresher messages contain more valuable information. This consensus is rooted in the understanding that fresh information ensures more informed and goal-oriented decision-making.
@@ II-C Mean Square Error (MSE)
MSE is a well-known metric that characterizes the accuracy aspect of information. Defined as Euclidean distance between transceivers’ statuses, MSE is used for timely reconstruction-oriented communications. In the MSE-oriented design, a status at the source is deemed significant if it considerably deviates from the estimated status at the monitor. When the sample policy is content-agnostic, the MSE can be expressed as a nonlinear function of AoI [14], making it a specialized form of VoI. MSE emphasizes the impact of the severity of E2E status mismatch on the effectiveness.
@@ II-E Age of Incorrect Information (AoII)
AoII is a new metric that orchestrates the strengths of MSE and AoS to measure relevance property [10]. As shown in Table I, AoII introduces three key innovations: First, AoII integrates AoS with MSE by multiplying the variants of them together, which incorporates both the duration and severity of E2E mismatch. Second, AoII incorporates a nonlinear penalty function predicated on AoS, drawing parallels to the penalty function in the Value of Information (VoI) relative to AoI. Third, AoII introduces a generalized error gap function, broadening the traditional Euclidean-distance-based MSE depictions. In this regard, AoII addresses that both severity and duration of E2E statues mismatch affect the effectiveness.
@@ II-G Urgency of Information (UoI)
While previous metrics focus solely on the relevance of the source, practical scenarios demonstrate that the surrounding environment also influences information relevance. For example, a self-driving car in complex situations like traffic jams would require more frequent status updates for safety, unlike when it’s on a clear highway. Addressing this, UoI is the first metric to link environment factors with information relevance [11]. Particularly, UoI introduces an environment-aware weighted coefficient based on the environment urgency level. The final UoI formula is a product of this coefficient and the error gap function, as detailed in Table I.
@@ III One More Step Forward: the Goal-oriented Tensor
The above metrics each examine effectiveness from varied indirect perspectives. In this section, we introduce a unified metric designed to gauge the relevance of information in a goal-oriented manner and seek to bring coherence to these various metrics.
@@ III-A Existing Metrics: Inherent Connection
From Section II and Table I, the existing metrics share two fundamental components: $i$) a content-aware error cost function $g(X(t),\hat{X}_{t})$ and $ii$) a content-independent weighted coefficient $\Phi(t)$. The former one describes the real-time E2E distance (like MSE or other error gap function) or mismatch cost (like Cost of Actuation Error). The latter one exerts a multiplicative effect on the error cost function. In this way, the inherent connection among the existing metrics is that they are all contingent upon a triple-tuple consisting of the status $X\left(t\right)$, the estimated status $\hat{X}\left(t\right)$, and the context $\Phi\left(t\right)$. This connection motivates us to propose a tensor-based approach to unify the existing metrics.
@@ III-C Constructing Generalized Goal-Oriented Tensor
The values in the generalized GoT manifest the cost of E2E statuses mismatch under a specific context status. The values in the GoT should relate to the goal directly. Here a detailed method to tailor a GoT based on the goal is given.
@@
$\bullet$ Step 2: Defining the semantic status set $\mathcal{S}$ and the context status set $\mathcal{V}$. Both sets are constructed as finite collections of discrete semantic segments, with the semantic status $X(t)\in\mathcal{S}$ and $\Phi(t)\in\mathcal{V}$ determined through semantic quantization (refer to Section IV). For example, $\mathcal{S}$ could involve semantic fragments, whereas $\mathcal{V}$ could encapsulate different weather conditions surrounding the observed source.
$\bullet$ Step 3: Determining the decision strategy in terms of $\hat{X}\left(t\right)$. Status update will be converted to a decision update at the actuators. This decision will turn back to affect the goal-oriented effectiveness of the system.
$\bullet$ Step 4: Evaluating the specific cost. There are three types of costs: $i$) the status inherent cost $\mathcal{C}_{1}(X(t),\Phi(t))$, which represents the cost under $X\left(t\right)$ and $\Phi\left(t\right)$ in the absence of external influences; $ii$) the decision gain $\mathcal{C}_{2}(\hat{X}(t))$, which is based on the fact that a good $\hat{X}(t)$-dependent decision could reduce the severity; iiiiii) the decision cost $\mathcal{C}_{3}(\hat{X}(t))$, which corresponds to the resource overhead due to a specific decision based on $\hat{X}(t)$
$\bullet$ Step 5: Calculating the GoT. The tensor value, given a specific triple-tuple $\langle X(t),\hat{X}(t),\Phi(t)\rangle$ and a determined decision mapping approach, is calculated by
$\mathcal{C}_{1}(X(t),\Phi(t))-\mathcal{C}_{3}(\hat{X}(t))]^{+}+\mathcal{C}_{2}(\hat{X}(t)),$ (1)
where the ramp function $\cdot$$^{+}$ ensures that any decision gain $\mathcal{C}_{3}$ will not reduce the cost below 0.
III-D How does GoT Address the Semantics and Effectiveness?
1) Semantics: The GoT addresses semantics from dual perspectives. First, it extends the distance of E2E meaningless raw data mismatch to the cost of context-specific E2E semantic mismatches. The interpretation of semantics as meaning is thus addressed. Second, the E2E semantic mismatch cost serves as a gauge of the relevance of information (or relevance of semantics as $X(t)$ represents semantic status). A higher context-specific mismatch cost indicates greater semantic relevance and vice versa, emphasizing the role of semantics in determining relevance.
2) Effectiveness: The tensor value within GoT is determined by both the E2E semantic mismatch and the decision strategy, representing the real-world cost due to these mismatches. Through GoT optimization, one can enhance decision-making and diminish significant E2E semantic inconsistencies. As a result, the long-term real-world cost is effectively reduced, highlighting GoT’s focus on effectiveness.
@@ IV A Holistic Framework for Goal-Oriented Semantic Communication
This section introduces a comprehensive goal-oriented semantic framework, depicted in Fig. 2, with the GoT acting as the central metric. The modules of our framework are elaborated as follows:
Semantic Quantization. Status update packets are generated when the sensors collect continuous raw data. The types of metadata are various, e.g., a piece of video/speech, a process of temperature variations, a fragment of moving track, or even their multi-modal combinations. Raw data are fed into a semantic quantizer for semantic extractions, producing semantics $X(t)$ and context $\Phi(t)$. This quantizer converts continuous, non-meaningful metadata into discrete semantic pieces. Leveraging DL-based methods, this transformation reduces the data volume per packet and facilitates the subsequent sparse semantics-aware sampling design.
@@
Goal-oriented Coding. Traditionally, channel coding strategies aim for perfect reconstructions. This method overlooks the delay-caused E2E status mismatch and its cost to the goal-oriented actuation. To address this issue, E2E decoding errors that may result in high costs can be assigned larger code distances. This allocation strategy reduces the probability of costly mismatches occurring. Conversely, decoding errors associated with lower costs can be allotted shorter code distances. By adopting this module, the system effectiveness can be further enhanced.